OpenStudy (anonymous):

what is the vertex form to:y=-1/2x^2+8/3x-25/3 ?

OpenStudy (anonymous):

y=-1/2x^2+8/3x-25/3 y=-1/2(x^2-(16/3)x+50/3) y=-1/2((x-8/3)^2+86/9) y=-1/2(x-8/3)^2) - 43/9